### v1

@@ 1 Introduction
The existing generation of atom interferometers have provided state-of-the-art measurements of accelerations Canuel:2006 ; Templier:2022 , rotations Gustavson:1997 ; Gustavson:2000 ; Durfee:2006 ; Gauguet:2009 ; Gautier:2022 , gravitational fields peters_measurement_1999 ; peters_high-precision_2001 ; altin_precision_2013 ; Hu:2013 ; Farah:2014 ; hardman_simultaneous_2016 ; Zhang:2023b , gravity gradients Snadden:1998 ; sorrentino_sensitivity_2014 ; Biedermann:2015 ; damico_bragg_2016 ; Asenbaum:2017 ; Janvier:2022 , the fine structure constant parker_measurement_2018 ; yu_atom-interferometry_2019 ; morel_determination_2020 , and Newton’s gravitational constant Rosi:2014 . With sufficient miniaturization and ruggedization, quantum sensors based on atom interferometry could enable new capabilities in navigation Jekeli:2005 ; Battelier:2016 ; Narducci:2022 ; wang_enhancing_2021 ; Wright:2022 ; Phillips:2022 , civil engineering risk management metje_seeing_2011 ; boddice_capability_2017 ; Stray:2022 , mineral exploration and recovery van_leeuwen_bhp_2000 ; bongs_taking_2019 , groundwater mapping and monitoring schilling_gravity_2020 , and geodesy Stockton2011 ; Migliaccio:2019 ; Trimeche:2019 ; Leveque:2021 . Atom interferometers are presently being developed for mobile operation on dynamic platforms, and have been deployed on ships Bidel:2018 ; Wu:2023 , aircraft Geiger2011 ; Bidel:2020 ; Bidel:2023 , and in microgravity environments onboard sounding rockets becker_space-borne_2018 ; Lachmann:2021 and the International Space Station williams_pathfinder_2024 . Spaceborne operation in particular has provided a strong motivation for next-generation atom-interferometric development, since it could progress key questions in fundamental physics through low-frequency-band gravitational wave detection dimopoulos_atomic_2008 , weak equivalence principle violation tests dimopoulos_testing_2007 ; williams_quantum_2016 , and novel experiments into dark energy burrage_using_2016 ; sabulsky_experiment_2019 , dark matter geraci_sensitivity_2016 ; badurina_aion_2020 , and quantum gravity Haine:2021 ; margalit_realization_2021 .
@@
There is a worldwide effort to decrease the size, weight, and power (SWaP) of atomic inertial sensors whilst maintaining sensitivity, accuracy, and stability on dynamic platforms in real-world environments fang_metrology_2016 ; bongs_taking_2019 ; Geiger:2020 ; Narducci:2022 . Efforts to address these challenges have largely focussed on improving the performance of the standard three-pulse Mach-Zehnder (MZ) atom interferometer, through innovations such as large momentum transfer atom optics Clade:2009 ; Muller:2008b ; Chiow:2011 ; McDonald:2013b ; Kotru:2015 ; Gebbe:2021 ; Wilkason:2022 ; Beguin:2023 , improved atomic source quality and production rate Debs:2011 ; robins_atom_2013 ; szigeti_why_2012 ; Deppner:2021 ; Hensel:2021 ; Chen:2022 ; Lee:2022 , novel state readout wigley_readout-delay-free_2019 ; Piccon:2022 ; benaicha:2024 , error-robust quantum control Saywell:2020 ; Saywell:2023 ; Saywell:2023a ; Wang:2024 ; Rodzinka:2024 , and overcoming the shot-noise limit through quantum entanglement generated from atom-atom esteve_squeezing_2008 ; appel_mesoscopic_2009 ; lucke_twin_2011 ; Haine:2011 ; hamley_spin-nematic_2012 ; lucke_detecting_2014 ; muessel_twist-and-turn_2015 ; lange_entanglement_2018 ; Szigeti:2020 ; Szigeti:2021 or atom-light hald_spin_1999 ; leroux_implementation_2010 ; schleier-smith_squeezing_2010 ; sewell_magnetic_2012 ; Haine:2013 ; Szigeti:2014b ; Haine:2016 ; Kritsotakis:2021 ; Fuderer:2023 interactions. However, another approach is to consider alternatives to the standard MZ atom interferometer, which could relax certain technological requirements and provide advantages in tight-SWaP situations.
@@
One alternative interferometry configuration is the atomic analogue of a Fabry-Perot interferometer (FPI). In an optical FPI, light enters a cavity formed by two parallel mirrors, and a resonant spectra is obtained by scanning the incident wavelength. Optical FPIs have been extensively used in a number of spectroscopic drever_laser_1983 ; deventer_comparison_1990 ; xue_pulsed_2016 and sensing yoshino_fiber-optic_1982 ; taylor_principles_1998 applications. In an atomic FPI, the incoming light is replaced by atomic matter-waves and the mirrors are replaced by laser-induced potential barriers. Previous theoretical studies into atomic FPIs have demonstrated that an ultracold Bose source can produce high contrast Fabry-Perot interference fringes carusotto_nonlinear_2001 ; paul_nonlinear_2005 ; paul_nonlinear_2007 ; rapedius_barrier_2008 ; ernst_transport_2010 , characterised their resonance properties dutt_smooth_2010 ; damon_reduction_2014 , and investigated the potential use of atomic FPIs in velocity selection wilkens_Fabry-Perot_1993 ; ruschhaupt_velocity_2005 and angle selection valagiannopoulos_quantum_2019 . For the experimental regimes achievable with current technology, a narrow momentum width source such as a Bose-Einstein condensate (BEC) is needed to achieve the high contrast resonant transmission peaks required for useful sensing manju_atomic_2020 . Operating in a regime where atom-atom collisional interactions are negligible is also highly desirable, and can be obtained through a Feshbach resonance Roberts:1998 ; Kuhn:2014 ; Everitt:2017 . The suitability of a noninteracting BEC source for atomic Fabry-Perot interferometry has been validated in a recent experimental demonstration using a 39K BEC source and optical barrier potentials formed using a digital micromirror device Eid:2024 .
@@
In this paper, we study the application of an atomic FPI as an acceleration sensor. A previous work has considered the suitability of atomic FPIs for gravimetry through a numerical simulation analysis schach_tunneling_2022 . Here we take a complementary analytic approach that aims to give deeper insights into the optimal parameter regime and performance limits of an atomic FPI acceleration sensor. This approach allows us to assess whether there are regimes where an atomic FPI could potentially offer superior performance as an acceleration sensor compared to MZ interferometry. We are particularly interested in situations where device size is highly constrained, such as in space-based applications. Since we are interested in fundamental performance limits, we consider only the case of a non-interacting BEC in this work.
Specifically, in Section 2 we derive an analytic expression for the transmission of a non-interacting BEC through the atomic FPI in an accelerating field, which we validate by numeric simulation of the Schrödinger equation for $N$ non-interacting particles. We first study in Section 3.1 the ideal case of an infinitely narrow momentum width source, and from the transmission derive an approximate expression for the optimum Fisher information (and consequently acceleration sensitivity). We then consider in Sections 3.3 and 3.4 atomic clouds with a finite momentum width, and study the effect of finite momentum width upon the free parameters which lead to optimal Fisher information. In each case, in Sections 3.2 and 3.5 respectively we compare the acceleration sensitivity of a space-based atomic FPI to a space-based MZ interferometer of equivalent device size, to assess the future potential of an atomic FPI as an alternative accelerometry device.
2 Methods
2.1 Model
@@
The probability of transmission through the double barrier system is given by the transmission coefficient xiao_resonant_2015
@@
In Figure 2a) we plot the transmission coefficient calculated via equation (3) (black solid curve) and numeric simulation (red dotted curve) corresponding to a range of $k_{i}/\kappa$ values, where $\kappa$ is a momentum length scale determined by $\kappa=\sqrt{2mV_{1}/\hbar^{2}}$. We use the cavity parameters for 85Rb determined in Ref. manju_atomic_2020 ($V_{1}=3.83\times 10^{-32}$ = $5.81\hbar\omega_{z}$, $w=1\upmu$m and $d=4\upmu$m) with trapping frequency $\omega_{z}=2\pi\times 10\text{ Hz}$ in the presence of an acceleration of $a=0$. Similarly, in Figure 2b) we plot the transmission coefficient for a fixed $k_{i}/\kappa=0.45$ and varying $a$. In both instances, we observe that the curve corresponding to equation (3) matches the simulation data very well, validating the analytic model.
2.2 Quantifying Acceleration Sensitivity
The smallest change in acceleration ($\delta a)$ detectable by an accelerometer quantifies the sensitivity of the device. For a cloud of $N$ non-interacting, uncorrelated atoms, this is given by the Cramér-Rao bound toth_quantum_2014
@@ 2.2 Quantifying Acceleration Sensitivity
where $F_{C}(a)$ is the per-particle classical Fisher information haine_mean-field_2016 ; Kritsotakis:2018 , given by
@@
3 Results and Discussion
3.1 Acceleration sensitivity of an atomic FPI with a plane matter-wave input
We begin our investigation into the sensitivity of an atomic FPI as an accelerometer by considering a BEC source with an infinitely narrow momentum width. This provides intuition for parameter dependencies of the sensitivity in the ideal case which optimises transmission through the FPI manju_atomic_2020 . $\tilde{F}_{C}$ varies with $\tilde{k}_{i}$ and we can estimate optimum $\tilde{k}_{i}$ that gives the maximum $\tilde{F}_{C}$ for each cavity length.
@@ 3.1 Acceleration sensitivity of an atomic FPI with a plane matter-wave input
The trend in Fig. 3 arises due to the changes in transmission peak properties with variation in the cavity length. As the cavity length increases, the linewidth of the resonant peaks gets narrower, leading to curves with higher slopes ($\partial T/\partial k$) manju_atomic_2020 . The relationship between the slope of the transmission spectra and acceleration sensitivity can be obtained as follows. Under uniform acceleration, the velocity of the cloud at the position of the first barrier is $v=v_{0}+at$, where $v_{0}$ is the initial velocity, $t$ is the time taken to reach the first barrier and $a$ is the acceleration. This yields
@@
3.2 Comparing Acceleration Sensitivities of an Atomic FPI and a MZ Interferometer
@@
In the case of a MZ interferometer, the per particle sensitivity is given by peters_high-precision_2001
@@
We can see immediately that making the device size $L_{t}$ smaller results in poorer sensitivity for both a MZ interferometer and an atomic FPI. However, equations (23b) and (26) show that the scaling of sensitivity with device length is different in each case. We are therefore motivated to study the variation of sensitivity $\delta a$ with length $L$, which we show for a range of finesse values $\mathcal{F}$ in Figure 5. Here optimum values are used for $k_{i},w$ and $d$ manju_atomic_2020 . For low length ($L<0.4\text{mm}$) and high finesse ($\mathcal{F}>100$), the fundamental sensitivity of an atomic FPI surpasses that of a MZ interferometer.
3.3 Space-Based Accelerometer Using a BEC With a Finite Momentum Width
Having compared an atomic FPI to a MZ interferometer for the ideal case of infinitely narrow momentum width, we now consider the more realistic case of finite momentum width. This immediately presents a new challenge: if this finite momentum width is greater than the cavity linewidth, only a portion of the atomic cloud (in momentum space) is on resonance and gets transmitted. This reduces the resonant transmission (i.e. $T_{\text{max}}<1$), so in order to achieve complete resonant transmission (i.e. $T_{\text{max}}=1$) we are immediately restricted to cavities with linewidth greater than the finite momentum width of the cloud. As linewidth decreases as cavity width increases, this puts an upper bound on the cavity width for an effective atomic FPI manju_atomic_2020 . For the cloud of finite momentum width, we consider a Gaussian of full width at half maximum (FWHM) of $\Delta k=0.008k_{0}$, where here $k_{0}=2\pi/(780\text{nm})$. Note that $\Delta k=2\sqrt{2\ln 2}/\sigma_{c}$, where $\sigma_{c}/\sqrt{2}$ is the standard deviation of the Gaussian density profile, or equivalently a Gaussian $k$-space density profile with standard deviation $1/(\sqrt{2}\sigma_{c})$.
@@ 3.3 Space-Based Accelerometer Using a BEC With a Finite Momentum Width
In Figure 6, we plot $\tilde{F}_{C}$ as a function $\tilde{k}_{i}$ for an atomic cloud with infinitely narrow and finite momentum width in Figure 6a) and Figure 6b) respectively. We use the parameters $w\kappa=1$, $w=1\upmu\text{m}$ and $\kappa=0.124k_{0}$, as determined in Ref. manju_atomic_2020 .
@@
3.4 Optimising Cavity Length and Corresponding Acceleration Sensitivity for a Particular Momentum Width
As before, we calculate the Fisher information $\tilde{F}_{C}$ as a function of initial momentum kick $\tilde{k}_{i}$ for a range of cavity widths $\tilde{d}$ and momentum widths $\Delta k=2\sqrt{2\ln 2}/\sigma_{c}$ which gives a maximum $\tilde{F}_{C}$. For each cavity length there is an optimum $\tilde{k}_{i}$ corresponding to the optimum Fisher information $\tilde{F}_{C}$ and this optimum point changes with cavity length. However, the optimum $\tilde{k}_{i}$ is not changed by variation in the momentum width $\Delta k$. This is because the optimum initial momentum kick $\tilde{k}_{i}$ depends upon the location of the first resonant peak, which is independent of the momentum width of the source. However, the momentum width $\Delta k$ does affect the height of the transmission peak, and therefore the optimum $\tilde{F}_{C}$ manju_atomic_2020 . We thus plot the maximum $\tilde{F}_{C}$ that occurs at the optimum $\tilde{k}_{i}$ as a function of cavity width $\tilde{d}$ in Figure 7 for different source momentum widths $\Delta k$.
Unlike in the infinitely narrow momentum width case (see Figure 3a)), there is not an unbounded increase in the optimum Fisher information $\tilde{F}_{C}$ with cavity width $\tilde{d}$. Rather, $\tilde{F}_{C}$ reaches a maximum and then decreases after the cavity width $\tilde{d}$ exceeds an optimal value. This optimal value arises because the cavity linewidth decreases as the cavity width increases manju_atomic_2020 . When the cavity width $\tilde{d}$ exceeds the optimal value, the corresponding linewidth becomes smaller than the momentum width $\Delta k$, and (as discussed earlier) transmission through the atomic FPI is suppressed and the optimal Fisher information $\tilde{F}_{C_{\text{opt}}}$ decreases.

### v2

@@ I Introduction
The existing generation of atom interferometers have provided state-of-the-art measurements of accelerations Canuel et al. (2006); Templier et al. (2022), rotations Gustavson et al. (1997, 2000); Durfee et al. (2006); Gauguet et al. (2009); Gautier et al. (2022), gravitational fields Peters et al. (1999, 2001); Altin et al. (2013); Hu et al. (2013); Farah et al. (2014); Hardman et al. (2016); Zhang et al. (2023), gravity gradients Snadden et al. (1998); Sorrentino et al. (2014); Biedermann et al. (2015); D’Amico et al. (2016); Asenbaum et al. (2017); Janvier et al. (2022), the fine structure constant Parker et al. (2018); Yu et al. (2019); Morel et al. (2020), and Newton’s gravitational constant Rosi et al. (2014). With sufficient miniaturization and ruggedization, quantum sensors based on atom interferometry could enable new capabilities in navigation Jekeli (2005); Battelier et al. (2016); Narducci et al. (2022); Wang et al. (2021); Wright et al. (2022); Phillips et al. (2022), civil engineering risk management Metje et al. (2011); Boddice et al. (2017); Stray et al. (2022), mineral exploration and recovery van Leeuwen (2000); Bongs et al. (2019), groundwater mapping and monitoring Schilling et al. (2020), and geodesy Stockton et al. (2011); Migliaccio et al. (2019); Trimeche et al. (2019); Lévèque et al. (2021). Atom interferometers are presently being developed for mobile operation on dynamic platforms, and have been deployed on ships Bidel et al. (2018); Wu et al. (2023), aircraft Geiger et al. (2011); Bidel et al. (2020, 2023), and in microgravity environments onboard sounding rockets Becker et al. (2018); Lachmann et al. (2021) and the International Space Station Williams et al. (2024). Spaceborne operation in particular has provided a strong motivation for next-generation atom-interferometric development, since it could progress key questions in fundamental physics through low-frequency-band gravitational wave detection Dimopoulos et al. (2008), weak equivalence principle violation tests Dimopoulos et al. (2007); Williams et al. (2016), and novel experiments into dark energy Burrage and Copeland (2016); Sabulsky et al. (2019), dark matter Geraci and Derevianko (2016); Badurina et al. (2020), and quantum gravity Haine (2021); Margalit et al. (2021).
@@
There is a worldwide effort to decrease the size, weight, and power (SWaP) of atomic inertial sensors whilst maintaining sensitivity, accuracy, and stability on dynamic platforms in real-world environments Fang et al. (2016); Bongs et al. (2019); Geiger et al. (2020); Narducci et al. (2022). Efforts to address these challenges have largely focussed on improving the performance of the standard three-pulse Mach-Zehnder (MZ) atom interferometer, through innovations such as large momentum transfer atom optics Cladé et al. (2009); Müller et al. (2008); Chiow et al. (2011); McDonald et al. (2013); Kotru et al. (2015); Gebbe et al. (2021); Wilkason et al. (2022); Béguin et al. (2023), improved atomic source quality and production rate Debs et al. (2011); Robins et al. (2013); Szigeti et al. (2012); Deppner et al. (2021); Hensel et al. (2021); Chen et al. (2022); Lee et al. (2022), novel state readout Wigley et al. (2019); Piccon et al. (2022); Ben-Aïcha et al. (2024), error-robust quantum control Saywell et al. (2020, 2023); Say (2023); Wang et al. (2024); Rodzinka et al. (2024), and overcoming the shot-noise limit through quantum entanglement generated from atom-atom Estève et al. (2008); Appel et al. (2009); Lücke et al. (2011); Haine and Ferris (2011); Hamley et al. (2012); Lücke et al. (2014); Muessel et al. (2015); Lange et al. (2018); Szigeti et al. (2020, 2021) or atom-light Hald et al. (1999); Leroux et al. (2010); Schleier-Smith et al. (2010); Sewell et al. (2012); Haine (2013); Szigeti et al. (2014); Haine and Lau (2016); Kritsotakis et al. (2021); Fuderer et al. (2023) interactions. However, another approach is to consider alternatives to the standard MZ atom interferometer, which could relax certain technological requirements and provide advantages in tight-SWaP situations.
@@
One alternative interferometry configuration is the atomic analogue of a Fabry-Perot interferometer (FPI). In an optical FPI, light enters a cavity formed by two parallel mirrors, and a resonant spectra is obtained by scanning the incident wavelength. Optical FPIs have been extensively used in a number of spectroscopic Drever et al. (1983); Deventer et al. (1990); Xue et al. (2016) and sensing Yoshino et al. (1982); Taylor (1998) applications. In an atomic FPI, the incoming light is replaced by atomic matter-waves and the mirrors are replaced by laser-induced potential barriers. Previous theoretical studies into atomic FPIs have demonstrated that an ultracold Bose source can produce high contrast Fabry-Perot interference fringes Carusotto (2001); Paul et al. (2005, 2007); Rapedius and Korsch (2008); Ernst et al. (2010), characterised their resonance properties Dutt and Kar (2010); Damon et al. (2014), and investigated the potential use of atomic FPIs in velocity selection Wilkens et al. (1993); Ruschhaupt et al. (2005) and angle selection Valagiannopoulos (2019). For the experimental regimes achievable with current technology, a narrow momentum width source such as a Bose-Einstein condensate (BEC) is needed to achieve the high contrast resonant transmission peaks required for useful sensing Manju et al. (2020). Operating in a regime where atom-atom collisional interactions are negligible is also highly desirable, and can be obtained through a Feshbach resonance Roberts et al. (1998); Kuhn et al. (2014); Everitt et al. (2017). The suitability of a noninteracting BEC source for atomic Fabry-Perot interferometry has been validated in a recent experimental demonstration using a 39K BEC source and optical barrier potentials formed using a digital micromirror device Eid et al. (2024).
@@
In this paper, we study the application of an atomic FPI as an acceleration sensor. A previous work has considered the suitability of atomic FPIs for gravimetry through a numerical simulation analysis Schach et al. (2022). Here we take a complementary analytic approach that aims to give deeper insights into the optimal parameter regime and performance limits of an atomic FPI acceleration sensor. This approach allows us to assess whether there are regimes where an atomic FPI could potentially offer superior performance as an acceleration sensor compared to MZ interferometry. We are particularly interested in situations where device size is highly constrained, such as in space-based applications. Since we are interested in fundamental performance limits, we consider only the case of a non-interacting BEC in this work.
Specifically, in Section II we derive an analytic expression for the transmission of a non-interacting BEC through the atomic FPI in an accelerating field, which we validate by numeric simulation of the Schrödinger equation for $N$ non-interacting particles. We first study in Section III.1 the ideal case of an infinitely narrow momentum width source, and from the transmission derive an approximate expression for the optimum Fisher information (and consequently acceleration sensitivity). We then consider in Sections III.3 and III.4 atomic clouds with a finite momentum width, and study the effect of finite momentum width upon the free parameters which lead to optimal Fisher information. In each case, in Sections III.2 and III.5 respectively we compare the acceleration sensitivity of a space-based atomic FPI to a space-based MZ interferometer of equivalent device size, to assess the future potential of an atomic FPI as an alternative accelerometry device.
II Methods
II.1 Model
@@
The probability of transmission through the double barrier system is given by the transmission coefficient Xiao et al. (2015)
@@
In Figure 2a) we plot the transmission coefficient calculated via equation (3) (black solid curve) and numeric simulation (red dotted curve) corresponding to a range of $k_{i}/\kappa$ values, where $\kappa$ is a momentum length scale determined by $\kappa=\sqrt{2mV_{1}/\hbar^{2}}$. We use the cavity parameters for 85Rb determined in Ref. Manju et al. (2020) ($V_{1}=3.83\times 10^{-32}$ = $5.81\hbar\omega_{z}$, $w=1\upmu$m and $d=4\upmu$m) with trapping frequency $\omega_{z}=2\pi\times 10\text{ Hz}$ in the presence of an acceleration of $a=0$. Similarly, in Figure 2b) we plot the transmission coefficient for a fixed $k_{i}/\kappa=0.45$ and varying $a$. In both instances, we observe that the curve corresponding to equation (3) matches the simulation data very well, validating the analytic model.
II.2 Quantifying Acceleration Sensitivity
The smallest change in acceleration ($\delta a)$ detectable by an accelerometer quantifies the sensitivity of the device. For a cloud of $N$ non-interacting, uncorrelated atoms, this is given by the Cramér-Rao bound Tóth and Apellaniz (2014)
@@ II.2 Quantifying Acceleration Sensitivity
where $F_{C}(a)$ is the per-particle classical Fisher information Haine (2016); Kritsotakis et al. (2018), given by
@@
III Results and Discussion
III.1 Acceleration sensitivity of an atomic FPI with a plane matter-wave input
We begin our investigation into the sensitivity of an atomic FPI as an accelerometer by considering a BEC source with an infinitely narrow momentum width. This provides intuition for parameter dependencies of the sensitivity in the ideal case which optimises transmission through the FPI Manju et al. (2020). $\tilde{F}_{C}$ varies with $\tilde{k}_{i}$ and we can estimate optimum $\tilde{k}_{i}$ that gives the maximum $\tilde{F}_{C}$ for each cavity length.
@@ III.1 Acceleration sensitivity of an atomic FPI with a plane matter-wave input
The trend in Fig. 3 arises due to the changes in transmission peak properties with variation in the cavity length. As the cavity length increases, the linewidth of the resonant peaks gets narrower, leading to curves with higher slopes ($\partial T/\partial k$) Manju et al. (2020). The relationship between the slope of the transmission spectra and acceleration sensitivity can be obtained as follows. Under uniform acceleration, the velocity of the cloud at the position of the first barrier is $v=v_{0}+at$, where $v_{0}$ is the initial velocity, $t$ is the time taken to reach the first barrier and $a$ is the acceleration. This yields
@@
III.2 Comparing Acceleration Sensitivities of an Atomic FPI and a MZ Interferometer
@@
In the case of a MZ interferometer, the per particle sensitivity is given by Peters et al. (2001)
@@
We can see immediately that making the device size $L_{t}$ smaller results in poorer sensitivity for both a MZ interferometer and an atomic FPI. However, equations (23b) and (26) show that the scaling of sensitivity with device length is different in each case. We are therefore motivated to study the variation of sensitivity $\delta a$ with length $L$, which we show for a range of finesse values $\mathcal{F}$ in Figure 5. Here optimum values are used for $k_{i},w$ and $d$ Manju et al. (2020). For low length ($L<0.4\text{mm}$) and high finesse ($\mathcal{F}>100$), the fundamental sensitivity of an atomic FPI surpasses that of a MZ interferometer.
III.3 Space-Based Accelerometer Using a BEC With a Finite Momentum Width
Having compared an atomic FPI to a MZ interferometer for the ideal case of infinitely narrow momentum width, we now consider the more realistic case of finite momentum width. This immediately presents a new challenge: if this finite momentum width is greater than the cavity linewidth, only a portion of the atomic cloud (in momentum space) is on resonance and gets transmitted. This reduces the resonant transmission (i.e. $T_{\text{max}}<1$), so in order to achieve complete resonant transmission (i.e. $T_{\text{max}}=1$) we are immediately restricted to cavities with linewidth greater than the finite momentum width of the cloud. As linewidth decreases as cavity width increases, this puts an upper bound on the cavity width for an effective atomic FPI Manju et al. (2020). For the cloud of finite momentum width, we consider a Gaussian of full width at half maximum (FWHM) of $\Delta k=0.008k_{0}$, where here $k_{0}=2\pi/(780\text{nm})$. Note that $\Delta k=2\sqrt{2\ln 2}/\sigma_{c}$, where $\sigma_{c}/\sqrt{2}$ is the standard deviation of the Gaussian density profile, or equivalently a Gaussian $k$-space density profile with standard deviation $1/(\sqrt{2}\sigma_{c})$.
@@ III.3 Space-Based Accelerometer Using a BEC With a Finite Momentum Width
In Figure 6, we plot $\tilde{F}_{C}$ as a function $\tilde{k}_{i}$ for an atomic cloud with infinitely narrow and finite momentum width in Figure 6a) and Figure 6b) respectively. We use the parameters $w\kappa=1$, $w=1\upmu\text{m}$ and $\kappa=0.124k_{0}$, as determined in Ref. Manju et al. (2020).
@@
III.4 Optimising Cavity Length and Corresponding Acceleration Sensitivity for a Particular Momentum Width
As before, we calculate the Fisher information $\tilde{F}_{C}$ as a function of initial momentum kick $\tilde{k}_{i}$ for a range of cavity widths $\tilde{d}$ and momentum widths $\Delta k=2\sqrt{2\ln 2}/\sigma_{c}$ which gives a maximum $\tilde{F}_{C}$. For each cavity length there is an optimum $\tilde{k}_{i}$ corresponding to the optimum Fisher information $\tilde{F}_{C}$ and this optimum point changes with cavity length. However, the optimum $\tilde{k}_{i}$ is not changed by variation in the momentum width $\Delta k$. This is because the optimum initial momentum kick $\tilde{k}_{i}$ depends upon the location of the first resonant peak, which is independent of the momentum width of the source. However, the momentum width $\Delta k$ does affect the height of the transmission peak, and therefore the optimum $\tilde{F}_{C}$ Manju et al. (2020). We thus plot the maximum $\tilde{F}_{C}$ that occurs at the optimum $\tilde{k}_{i}$ as a function of cavity width $\tilde{d}$ in Figure 7 for different source momentum widths $\Delta k$.
Unlike in the infinitely narrow momentum width case (see Figure 3a)), there is not an unbounded increase in the optimum Fisher information $\tilde{F}_{C}$ with cavity width $\tilde{d}$. Rather, $\tilde{F}_{C}$ reaches a maximum and then decreases after the cavity width $\tilde{d}$ exceeds an optimal value. This optimal value arises because the cavity linewidth decreases as the cavity width increases Manju et al. (2020). When the cavity width $\tilde{d}$ exceeds the optimal value, the corresponding linewidth becomes smaller than the momentum width $\Delta k$, and (as discussed earlier) transmission through the atomic FPI is suppressed and the optimal Fisher information $\tilde{F}_{C_{\text{opt}}}$ decreases.